OpenStudy (anonymous):

Which point is represented on the line represented by the equation y+4 =-5(x-3)? A. (-4,-5) B. (-5,-4) C. (3,-4) D. (-3,4)

OpenStudy (serenity74):

Make x = 0 and solve for y Then make y = 0 and solve for X. and then u get the coordinates.

OpenStudy (anonymous):

oh ok thnx.

OpenStudy (serenity74):

no problem :)

OpenStudy (serenity74):

did u get it?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

could u explain some more?

OpenStudy (serenity74):

Ohhh you didn't simplify it! That's why!

OpenStudy (anonymous):

I don't even think I got it correct.

OpenStudy (anonymous):

y+4=-5(0-3) y+4=15 is what i got but i dont know what to do after that

OpenStudy (serenity74):

Distribute it out and then make x=0 and y=0. I will give you an example. y+3 = 2(x-1) (distribute it) y+3=2x-2 (you then isolate the y) y = 2x-2+3 (then add up like terms) y = 2x +1 (after you get this in order to get the y value u make x=0) y = 2(0) + 1 = 1 y = 1 (in order to get the x value u make y = 0) 0 = 2x +1 -1 = 2x (then divide the two on both sides in order to isolate the x -1/2 = x and bam you get (-1/2, 1)

OpenStudy (serenity74):

Do u get it now?

OpenStudy (anonymous):

um its y+4 not y+3

OpenStudy (serenity74):

IK lol I gave you an example lol

OpenStudy (anonymous):

oh lol sorry

OpenStudy (anonymous):

okay let me try to do it.

OpenStudy (anonymous):

on the second line where u put y+3=2x-2 is that correct or is the 2 supposed to be a one?

OpenStudy (serenity74):

Do u get it now?

OpenStudy (anonymous):

sorry I still don't get it.